# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ 6.3.2 Relaxation timescales with relaxation paths
The large majority of unstable misalignments return to their original stability regime with co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths making up $\approx 76\%$ and $\approx 4\%$ of our sample respectively (see Table 3). Co $\rightarrow$ counter relaxations make up $\approx 15.8\%$, while we find only 71 instances ($\approx 2.3\%$) of counter $\rightarrow$ co relaxations. Given that only $\approx 6\%$ of galaxies in our sample at a given snipshot are counter-rotating, it is unsurprising that we find significantly fewer unstable misalignments originating from the counter-rotating stable regime.
Figure 7: Unstable misalignment relaxation pathways between $10<z<$ 1, showing co-rotating to co-rotating relaxations (upper left), co-rotating to counter-rotating relaxations (upper right), counter-rotating to counter-rotating relaxations (lower left), and counter-rotating to co-rotating relaxations (lower right). Relaxation paths are stacked at the first snipshot in the unstable regime $[160^{\circ}-20^{\circ}]$ (vertical grey line). Unstable misalignments will eventually relax back into the aligned ($<30^{\circ}$) or counter-rotating ($>150^{\circ}$) regimes (light grey shaded regions). Galaxies are assumed to have relaxed upon spending $>0.1$ Gyr in a consecutive stable regime of $<20^{\circ}$ or $>160^{\circ}$ (dark grey shaded regions). Median relaxation paths (solid black line) and $1\sigma$ percentile relaxation paths (dashed black lines) are shown for each relaxation type. Longer relaxation times have had their hue darkened to stand out, while short relaxation times have increased transparency.
@@
We see a clear positive relationship between the relaxation time and the displaced angle up to $\approx 90^{\circ}$, showing that unstable misalignments displaced by a larger amount from the nearest stability regime will take longer to relax. We also see a population of unstable misalignments that return to their original stability regime despite being displaced angles of $>90^{\circ}$. To first order, unstable misalignments are expected to relax toward the nearest stable regime. This population of unstable misalignments displaced by $>90^{\circ}$ is, therefore, indicative of additional processes acting on the gas (e.g. smooth misaligned accretion) that prevent it from relaxing toward the nearest dynamically stable regime. Unstable misalignments displaying this behaviour tend to show significant enhancements to their relaxation timescales with $\gtrsim 3\bar{t}_{\rm{torque}}$.
@@
We see from Figure 7 that co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths show similar behaviour and the majority are displaced by only a small angle from the stable co- and counter-rotating regimes, respectively. As small displacements require less angular momentum dissipation, shorter relaxation timescales are expected for the average co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths in our sample. For instance, the large majority ($\approx 73\%$) of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}<1$ relaxations are displaced by an angle of $<40^{\circ}$ from coplanarity.
For relaxations that transition between stable regimes, this automatically involves displaced angles of $>90^{\circ}$. As more angular momentum must be dissipated in order to relax into one of the stable regimes, co $\rightarrow$ counter and counter $\rightarrow$ co relaxation paths are expected to return longer relaxation timescales. We also see that co $\rightarrow$ co and counter $\rightarrow$ counter relaxations follow similar paths. This makes sense as eagle models stellar particles as non-collisional. Thus, to first order, co-rotating and counter-rotating systems can be considered dynamically-equivalent from the perspective of a misaligned gas particle in a polar orbit.
@@ 6.3.3 Morphological dependence on relaxation timescales
To test for significance between these distributions, we run a two-sample KS test between LTG and ETG $t_{\rm{relax}}$ distributions. We find significant results for $t_{\rm{relax}}$ (KS-test statistic $=0.22$, p-value $=7.6\times 10^{-25}$), with a decreasing significance for $t_{\rm{relax}}/\bar{t}_{\rm{dyn}}$ (KS-test statistic $=0.15$, p-value $=2.3\times 10^{-11}$). We find no significant difference between morphologies in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (KS-test statistic $=0.03$, p-value $=0.85$). This likely encapsulates the greater ability of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ to account for dynamical and morphological properties of galaxies. We find no changes in the significance of the results using a stricter morphological classification of $\bar{\kappa}_{\rm{co}}^{\rm{*}}<0.35$ for ETGs and $\bar{\kappa}_{\rm{co}}^{\rm{*}}>0.45$ for LTGs.
@@
Interestingly, while the shape of the $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ distributions and the gradient of the log-linear relationship describing the distribution is similar between morphologies, the composition of relaxation paths is different (as seen clearly in Figure 9, with compositional fractions in Table 3). For instance, co $\rightarrow$ counter relaxations constitute $\approx 33\%$ of detected LTG misalignments, compared to only $\approx 16\%$ in ETGs. The longer inherent relaxation times of co $\rightarrow$ counter relaxation paths (as highlighted in Section 6.3.2) dominate the tail-end of the distributions for LTGs. This tends to increase the overall median relaxation time of LTGs ($0.38$ Gyr).
@@
A physical explanation may lie in the underlying differences in size and mass of the gasSF disc between morphologies. Given LTGs typically host larger star-forming gas discs, this sets a larger gas accretion threshold for the formation of unstable misalignments. This likely lowers the relative number of co $\rightarrow$ co misalignments that are otherwise common among ETGs. This could suggest larger stochastic accretion events such as gas-rich mergers or accretion through galaxy interactions are be responsible for many of these misalignments in LTGs.
LTGs and gas-rich galaxies are also found to preferentially reside in field and group environments (Dressler, 1980). Compared to virialized clusters (which are dominated by ETGs and gas-poor galaxies), galaxy interaction rates are expected to be higher in group environments (with the exception of the central massive galaxy). Likewise, the larger gas discs found in many LTGs create larger collisional cross-sections for gas stripping and mass exchange during close encounters with nearby gas-rich satellites. These encounters can significantly perturb the outskirts of the gas disc, which can manifest as an observable stellar-gas misalignment (e.g. Lu et al., 2021; Khim et al., 2021). As gas accretion from galaxy encounters can be more stochastic and violent (compared to the smooth accretion of e.g. halo cooling), these encounters may have a greater likelihood to flip the star-forming gas disc into a counter-rotating state, which disproportionately affects LTGs. This may hint at an underlying difference in drivers of misalignments between morphologies depending on the large-scale environment.
@@ 6.3.4 Relaxation paths and morphology
In Figure 10 we compare the relaxation times of individual relaxation paths between morphologies. Considering relaxation paths separately, we see a trend toward longer median relaxation times in ETGs compared to LTGs (see Table 2). For the dominant co $\rightarrow$ co relaxation paths in our sample, ETGs take $\approx 1.5$ times longer to relax compared to LTGs. Testing for significance between morphologies in terms of $\bar{t}_{\rm{relax}}$ returns significant results for each path; co $\rightarrow$ co (KS-test statistic $=0.25$, p-value $=2.2\times 10^{-21}$), counter $\rightarrow$ counter (KS-test statistic $=0.29$, p-value $=2.3\times 10^{-2}$), co $\rightarrow$ counter (KS-test statistic $=0.30$, p-value $=9.1\times 10^{-8}$), and counter $\rightarrow$ co (KS-test statistic $=0.54$, p-value $=1.3\times 10^{-2}$).
While we found no significant difference between morphologies in the overall normalised relaxation time distributions in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (see Section 6.3.3), we do find a difference when co $\rightarrow$ co relaxation paths are considered separately. For co $\rightarrow$ co relaxation paths, ETGs experience $\approx 1.2$ times longer normalised relaxation times compared to LTGs and this result is significant (KS-test statistic $=0.11$, p-value $=1.5\times 10^{-4}$). Median co $\rightarrow$ counter relaxation paths in ETGs are also $\approx 1.2$ longer compared to LTGs in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$. However, this result narrowly falls outside of being significant (KS-test statistic $=0.13$, p-value $=6.2\times 10^{-2}$). The remaining relaxation paths do not show any differences between ETGs and LTGs that return significant KS-test results.
@@ 6.3.5 Gas fraction dependence on relaxation timescales
In Figure 11 we show the variation in normalised relaxation time for a range of gasSF fractions within $<r_{50}$, independent of morphology. Henceforth we exclude short-duration unstable misalignments close to the resolution limit (i.e. $t_{\mathrm{relax}}<0.2$ Gyr), due to the incompleteness and the larger uncertainties associated with relaxation timescales in this range. We retain a sample of 2614 relaxations, of which 466 lie within $0.10<f_{\rm{gas,SF}}<$ 0.1, 1646 within $0.30.1<f_{\rm{gas,SF}}<$ 0.3, and 502 within $10.3<f_{\rm{gas,SF}}<$ 1.
We see a trend toward longer-duration unstable misalignments for lower $f_{\rm{gas,SF}}$. This is seen clearly between the middle and highest $f_{\rm{gas,SF}}$ sub-samples (two-sample KS-test statistic $=0.13$ and p-value $=4.6\times 10^{-5}$), and less strongly for the lowest and middle $f_{\rm{gas,SF}}$ sub-samples (two-sample KS-test statistic $=0.09$ and p-value $=5.5\times 10^{-3}$). This is also reflected in the median normalised relaxation times, with medians of $2.0t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, $1.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, and $1.4t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ for increasing $f_{\rm{gas,SF}}$ sub-samples. These results are unchanged when considering only co $\rightarrow$ co and counter $\rightarrow$ counter relaxations, and we conclude that bias from different relaxation path compositions is negligible. We expect these results to correlate with morphology because ETGs in eagle typically host lower star-forming gas fractions compared to LTGs (e.g. Casanueva et al., 2022), echoing existing observational results (see e.g. Ruffa et al., 2019a, b).
@@ 6.3.6 Central vs satellite galaxy dependence on relaxation timescales
We find a small trend toward longer normalised relaxation times in central galaxies compared to satellite galaxies with median relaxation times of $1.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ and $1.6t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, respectively (not shown). Comparing the distributions of normalised relaxation times between central and satellite galaxies we find this difference is significant (two-sample KS-test statistic $=0.08$ and p-value $=2.0\times 10^{-3}$).
@@
6.3.7 Ongoing gas inflow rate dependence on relaxation timescales
We estimate the ongoing gas inflow rate by tracing gas particles entering a sphere with a radius of $2r_{50}$. This allows us to consider gas accretion onto gas discs that extend beyond our kinematic aperture ($r_{50}$). From visual inspection, we ignore the gas inflow over the first 0.3 Gyr post-formation, which sets a lower limit of $t_{\mathrm{relax}}>0.3$ Gyr. This allows us to largely ignore the initial gas accretion that may have formed the misalignment and focus on ongoing gas accretion. We time-average the gas inflow rate, $\dot{M}_{\mathrm{gas}}$, for the remainder of the relaxation. We limit our analysis to central galaxies because a significant proportion of unstable misalignments in satellite galaxies are likely affected by gas-stripping (e.g. ram-pressure stripping; Khim et al., 2021) rather than gas accretion. We use 1148 relaxations for which the central galaxy classification remains unchanged for the duration of $t_{\rm{relax}}$ and for which the ongoing gas inflow rates can be estimated. Of these, 286 have ongoing gas accretion rates of $\dot{M}_{\mathrm{gas}}<1$ M yr-1, 773 lie within $101<\dot{M}_{\mathrm{gas}}<$ 10 M yr-1, and 89 have higher ongoing accretion rates of $\dot{M}_{\mathrm{gas}}>10$ M yr-1.
Figure 12: Normalised relaxation time distributions of central galaxies with gas inflow rates of $\dot{M}_{\mathrm{gas}}<1$ M yr-1 and $\dot{M}_{\mathrm{gas}}>10$ M yr-1, averaged 0.3 Gyr post-formation at $2r_{50}$. Percentages are given with respect to total relaxations within the inflow rate sub-sample. Errors are given as Poisson uncertainties. We find high gas inflow leads to longer normalised relaxation times than low inflow, with medians of $2.0t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and $2.8t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, respectively.
In Figure 12 we show how the normalised relaxation time varies with the time-averaged ongoing gas inflow rate (where we have omitted the intermediate inflow range for clarity). Central galaxies with higher gas inflow rates show longer-lived unstable misalignments. We find median normalised relaxation times of $2.0t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, $2.5t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, and $2.8t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ for $\dot{M}_{\mathrm{gas}}<1$ M yr-1 to $101<\dot{M}_{\mathrm{gas}}<$ 10 M yr-1 to $\dot{M}_{\mathrm{gas}}>10$ M yr-1, respectively. This effect is seen most clearly in the range $2\gtrsim t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}\gtrsim 8$. The difference between the lower and higher inflow galaxies result is significant (two-sample KS-test statistic $=0.16$ and p-value $=1.6\times 10^{-4}$). Further splitting the sample among ETGs and LTGs reveals the significance of this result to be driven primarily by central ETGs rather than central LTGs (not shown). Averaging the gas inflow rate after a longer time since initial misalignment formation (e.g. $>0.5$ Gyr) strengthens the dichotomy between low- and high-inflow unstable misalignments.
6.3.8 Environmental dependence on relaxation timescales
Finally, smooth accretion from misaligned halo cooling has been suggested as both a formation path of misalignments (e.g. Lagos et al., 2015), and a means to extend the lifetimes of an existing unstable misalignment (e.g. Bryant et al., 2019). Thus, we consider the effects of overall environment density on the relaxation distribution by using the halo mass, $M_{\rm{200}}$, as a proxy. These results are shown in Figure 13.
We limit ourselves to central galaxies for this analysis as gas cooling from a hot halo will primarily accrete onto the central galaxy (e.g. Fabian et al., 1984; Negri et al., 2014; Gaspari et al., 2017). We consider two mass ranges broadly corresponding to low-mass halos ($M_{\rm{200}}<10^{12.5}$ M) and intermediate to high-mass halos ($M_{\rm{200}}>10^{12.5}$ M). This provides a sample of 1699 relaxations with sub-sample sizes of 1615 (low-mass) and 84 (intermediate/high-mass). Because we find only 2 instances of central cluster galaxies (assumed to be $M_{\rm{200}}>10^{14}$ M), we only distinguish between low- and intermediate-mass environments.

### v2

@@ 6.3.2 Relaxation timescales with relaxation paths
The large majority of unstable misalignments return to their original stability regime with co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths making up $\approx 76$ per cent and $\approx 4$ per cent of our sample respectively (see Table 3). Co $\rightarrow$ counter relaxations make up $\approx 15.8$ per cent, while we find only 71 instances ($\approx 2.3$ per cent) of counter $\rightarrow$ co relaxations. Given that only $\approx 6$ per cent of galaxies in our sample at a given snipshot are counter-rotating, it is unsurprising that we find significantly fewer unstable misalignments originating from the counter-rotating stable regime.
Figure 8: Unstable misalignment relaxation pathways between $0<z<1$, showing co-rotating to co-rotating relaxations (upper left), co-rotating to counter-rotating relaxations (upper right), counter-rotating to counter-rotating relaxations (lower left), and counter-rotating to co-rotating relaxations (lower right). Relaxation paths are stacked at the first snipshot in the unstable regime $20^{\circ}-160^{\circ}]$ (vertical grey line). Unstable misalignments will eventually relax back into the aligned ($30^{\circ}$) or counter-rotating ($150^{\circ}$) regimes (light grey shaded regions). Galaxies are assumed to have relaxed upon spending >0.1>0.1 Gyr in a consecutive stable regime of $20^{\circ}$ or $160^{\circ}$ (dark grey shaded regions). Median relaxation paths (solid black line) and $1\sigma$ percentile relaxation paths (dashed black lines) are shown for each relaxation type. Longer relaxation times have had their hue darkened to stand out, while short relaxation times have increased transparency.
@@
We see a clear positive relationship between the relaxation time and the displaced angle up to $\approx 90^{\circ}$, showing that unstable misalignments displaced by a larger amount from the nearest stability regime will take longer to relax. We also see a population of unstable misalignments that return to their original stability regime despite being displaced angles of $90^{\circ}$. To first order, unstable misalignments are expected to relax toward the nearest stable regime. This population of unstable misalignments displaced by $90^{\circ}$ is, therefore, indicative of additional processes acting on the gas (e.g. smooth misaligned accretion) that prevent it from relaxing toward the nearest dynamically stable regime. Unstable misalignments displaying this behaviour tend to show significant enhancements to their relaxation timescales with $\gtrsim 3\bar{t}_{\rm{torque}}$.
@@
We see from Figure 8 that co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths show similar behaviour and the majority are displaced by only a small angle from the stable co- and counter-rotating regimes, respectively. As small displacements require less angular momentum dissipation, shorter relaxation timescales are expected for the average co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths in our sample. For instance, the large majority ($\approx 73$ per cent) of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}<1$ relaxations are displaced by an angle of $40^{\circ}$ from coplanarity.
For relaxations that transition between stable regimes, this automatically involves displaced angles of $90^{\circ}$. As more angular momentum must be dissipated in order to relax into one of the stable regimes, co $\rightarrow$ counter and counter $\rightarrow$ co relaxation paths are expected to return longer relaxation timescales. We also see that co $\rightarrow$ co and counter $\rightarrow$ counter relaxations follow similar paths. This makes sense as eagle models stellar particles as non-collisional. Thus, to first order, co-rotating and counter-rotating systems can be considered dynamically-equivalent from the perspective of a misaligned gas particle in a polar orbit.
@@ 6.3.3 Morphological dependence on relaxation timescales
To test for significance between these distributions, we run a two-sample KS test between LTG and ETG $t_{\rm{relax}}$ distributions. We find significant results for $t_{\rm{relax}}$ (KS-test statistic =0.22=0.22, p-value $7.6\times 10^{-25}$), with a decreasing significance for $t_{\rm{relax}}/\bar{t}_{\rm{dyn}}$ (KS-test statistic =0.15=0.15, p-value $2.3\times 10^{-11}$). We find no significant difference between morphologies in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (KS-test statistic =0.03=0.03, p-value =0.85=0.85). This likely encapsulates the greater ability of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ to account for dynamical and morphological properties of galaxies. We find no changes in the significance of the results using a stricter morphological classification of $\bar{\kappa}_{\rm{co}}^{\rm{*}}<0.35$ for ETGs and $\bar{\kappa}_{\rm{co}}^{\rm{*}}>0.45$ for LTGs.
@@
Interestingly, while the shape of the $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ distributions and the gradient of the log-linear relationship describing the distribution is similar between morphologies, the composition of relaxation paths is different (as seen clearly in Figure 10, with compositional fractions in Table 3). For instance, co $\rightarrow$ counter relaxations constitute $\approx 33$ per cent of detected LTG misalignments, compared to only $\approx 16$ per cent in ETGs. The longer inherent relaxation times of co $\rightarrow$ counter relaxation paths (as highlighted in Section 6.3.2) dominate the tail-end of the distributions for LTGs. This tends to increase the overall median relaxation time of LTGs ($0.38$ Gyr).
@@
A physical explanation may lie in the underlying differences in size and mass of the gasSF disc between morphologies. Given LTGs typically host larger star-forming gas discs, this sets a larger gas accretion threshold for the formation of unstable misalignments. This likely lowers the relative number of co $\rightarrow$ co misalignments that are otherwise common among ETGs. This could suggest larger stochastic accretion events such as gas-rich mergers or accretion through galaxy interactions are responsible for many of these misalignments in LTGs.
Furthermore, eagle reproduces the observed galaxy morphology-density relation (Dressler, 1980), with LTGs found to preferentially reside in lower density environments (Pfeffer et al., 2023). Compared to virialized high-density environments (which are dominated by gas-poor ETGs), galaxy interaction rates tend to be higher in lower-density environments (Ghigna et al., 1998; Pfeffer et al., 2023). Likewise, the larger gas discs found in many LTGs create larger collisional cross-sections for gas stripping and mass exchange during close encounters with nearby gas-rich satellites. These encounters can significantly perturb the outskirts of the gas disc, which can manifest as an observable stellar-gas misalignment (e.g. Lu et al., 2021; Khim et al., 2021). As gas accretion from galaxy encounters can be more stochastic and violent (compared to the smooth accretion of e.g. halo cooling), these encounters may have a greater likelihood to flip the star-forming gas disc into a counter-rotating state, which disproportionately affects LTGs. This may hint at an underlying difference in drivers of misalignments between morphologies depending on the large-scale environment.
@@ 6.3.4 Relaxation paths and morphology
In Figure 11 we compare the relaxation times of individual relaxation paths between morphologies. Considering relaxation paths separately, we see a trend toward longer median relaxation times in ETGs compared to LTGs (see Table 2). For the dominant co $\rightarrow$ co relaxation paths in our sample, ETGs take $\approx 1.5$ times longer to relax compared to LTGs. Testing for significance between morphologies in terms of $\bar{t}_{\rm{relax}}$ returns significant results for each path; co $\rightarrow$ co (KS-test statistic =0.25=0.25, p-value $2.2\times 10^{-21}$), counter $\rightarrow$ counter (KS-test statistic =0.29=0.29, p-value $2.3\times 10^{-2}$), co $\rightarrow$ counter (KS-test statistic =0.30=0.30, p-value $9.1\times 10^{-8}$), and counter $\rightarrow$ co (KS-test statistic =0.54=0.54, p-value $1.3\times 10^{-2}$).
While we found no significant difference between morphologies in the overall normalised relaxation time distributions in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (see Section 6.3.3), we do find a difference when co $\rightarrow$ co relaxation paths are considered separately. For co $\rightarrow$ co relaxation paths, ETGs experience $\approx 1.2$ times longer normalised relaxation times compared to LTGs and this result is significant (KS-test statistic =0.11=0.11, p-value $1.5\times 10^{-4}$). Median co $\rightarrow$ counter relaxation paths in ETGs are also $\approx 1.2$ longer compared to LTGs in terms of $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$. However, this result narrowly falls outside of being significant (KS-test statistic =0.13=0.13, p-value $6.2\times 10^{-2}$). The remaining relaxation paths do not show any differences between ETGs and LTGs that return significant KS-test results.
@@ 6.3.5 Gas fraction dependence on relaxation timescales
In Figure 12 we show the variation in normalised relaxation time for a range of gasSF fractions within $r_{50}$, independent of morphology. Henceforth we exclude short-duration unstable misalignments close to the resolution limit (i.e. $t_{\mathrm{relax}}<0.2$ Gyr), due to the incompleteness and the larger uncertainties associated with relaxation timescales in this range. We retain a sample of 2614 relaxations, of which 466 lie within $0<f_{\rm{gas,SF}}<0.1$, 1646 within $0.1<f_{\rm{gas,SF}}<0.3$, and 502 within $0.3<f_{\rm{gas,SF}}<1$.
We see a trend toward longer-duration unstable misalignments for lower $f_{\rm{gas,SF}}$. This is seen clearly between the middle and highest $f_{\rm{gas,SF}}$ sub-samples (two-sample KS-test statistic =0.13=0.13 and p-value $4.6\times 10^{-5}$), and less strongly for the lowest and middle $f_{\rm{gas,SF}}$ sub-samples (two-sample KS-test statistic =0.09=0.09 and p-value $5.5\times 10^{-3}$). This is also reflected in the median normalised relaxation times, with medians of $2.0t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, $1.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, and $1.4t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ for increasing $f_{\rm{gas,SF}}$ sub-samples. These results are unchanged when considering only co $\rightarrow$ co and counter $\rightarrow$ counter relaxations, and we conclude that bias from different relaxation path compositions is negligible. We expect these results to correlate with morphology because ETGs in eagle typically host lower star-forming gas fractions compared to LTGs (e.g. Casanueva et al., 2022), echoing existing observational results (see e.g. Ruffa et al., 2019a, b).
@@
6.3.6 Ongoing gas inflow rate dependence on relaxation timescales
We estimate the ongoing gas inflow rate by tracing gas particles entering a sphere with a radius of $2r_{50}$. This allows us to consider gas accretion onto gas discs that extend beyond our kinematic aperture ($r_{50}$). From visual inspection, we ignore the gas inflow over the first 0.3 Gyr post-formation, which sets a lower limit of $t_{\mathrm{relax}}>0.3$ Gyr. This allows us to largely ignore the initial gas accretion that may have formed the misalignment and focus on ongoing gas accretion. We time-average the gas inflow rate, $\dot{M}_{\mathrm{gas}}$, for the remainder of the relaxation. We limit our analysis to central galaxies because a significant proportion of unstable misalignments in satellite galaxies are likely affected by gas-stripping (e.g. ram-pressure stripping; Khim et al., 2021) rather than gas accretion. We use 1148 relaxations for which the central galaxy classification remains unchanged for the duration of $t_{\rm{relax}}$ and for which the ongoing gas inflow rates can be estimated. Of these, 286 have ongoing gas accretion rates of $\dot{M}_{\mathrm{gas}}<1$ M yr-1, 773 lie within $1<\dot{M}_{\mathrm{gas}}<10$ M yr-1, and 89 have higher ongoing accretion rates of $\dot{M}_{\mathrm{gas}}>10$ M yr-1.
Figure 13: Normalised relaxation time distributions of central galaxies with gas inflow rates of $\dot{M}_{\mathrm{gas}}<1$ M yr-1 and $\dot{M}_{\mathrm{gas}}>10$ M yr-1, averaged 0.3 Gyr post-formation at $2r_{50}$. Percentages are given with respect to total relaxations within the inflow rate sub-sample. Errors are given as Poisson uncertainties. We find high gas inflow leads to longer normalised relaxation times than low inflow, with medians of $2.8t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and $2.0t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, respectively.
In Figure 13 we show how the normalised relaxation time varies with the time-averaged ongoing gas inflow rate (where we have omitted the intermediate inflow range for clarity). Central galaxies with higher gas inflow rates show longer-lived unstable misalignments. We find median normalised relaxation times of $2.0t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, $2.5t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$, and $2.8t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ for $\dot{M}_{\mathrm{gas}}<1$ M yr-1 to $1<\dot{M}_{\mathrm{gas}}<10$ M yr-1 to $\dot{M}_{\mathrm{gas}}>10$ M yr-1, respectively. This effect is seen most clearly in the range $2\gtrsim t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}\gtrsim 8$. The difference between the lower and higher inflow galaxies result is significant (two-sample KS-test statistic =0.16=0.16 and p-value $1.6\times 10^{-4}$). Further splitting the sample among ETGs and LTGs reveals the significance of this result to be driven primarily by central ETGs rather than central LTGs (not shown). Averaging the gas inflow rate after a longer time since initial misalignment formation (e.g. >0.5>0.5 Gyr) strengthens the dichotomy between low- and high-inflow unstable misalignments.
@@ 6.3.7 Central vs satellite galaxy dependence on relaxation timescales
We find a small trend toward longer normalised relaxation times in central galaxies compared to satellite galaxies with median relaxation times of $1.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ and $1.6t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, respectively (not shown). Comparing the distributions of normalised relaxation times between central and satellite galaxies we find this difference is significant (two-sample KS-test statistic =0.08=0.08 and p-value $2.0\times 10^{-3}$).
@@
6.3.8 Halo mass dependence on relaxation timescales
Finally, smooth accretion from misaligned halo cooling has been suggested as both a formation path of misalignments (e.g. Lagos et al., 2015), and a means to extend the lifetimes of an existing unstable misalignment (e.g. Bryant et al., 2019). Thus, we consider the effects of overall environment density (and by proxy, larger halo gas reservoirs) on the relaxation distribution by using the halo mass, $M_{\rm{200c}}$, as a proxy. These results are shown in Figure 14.
We limit ourselves to central galaxies for this analysis as gas cooling from a hot halo will primarily accrete onto the central galaxy (e.g. Fabian et al., 1984; Negri et al., 2014; Gaspari et al., 2017). We consider two mass ranges broadly corresponding to low-mass halos ($M_{\rm{200c}}<10^{12.5}$ M) and intermediate to high-mass halos ($M_{\rm{200c}}>10^{12.5}$ M). This provides a sample of 1699 relaxations with sub-sample sizes of 1615 (low-mass) and 84 (intermediate/high-mass). Because we find only 2 instances of central cluster galaxies (assumed to be $M_{\rm{200c}}>10^{14}$ M) among our sample, we only distinguish between low- and intermediate-mass environments.